# Variational Zero-shot Multispectral Pansharpening

### v1

@@ III-C Optimization for the PSDip model
PSDip can be conveniently solved by iteratively updating. Let $\mathcal{L}(\mathcal{X},\theta)$ denote the objective function of problem (7)
@@
Then, $\mathcal{X}$ can be updated by
@@
We find that the initialization of $\theta$ has a significant impact on the above iterative updating method. Note that although $\mathcal{G}$ could be regularized by the network structure to some extent, its value can still be easily influenced by outside guidance, e.g. $\mathcal{X}_{t}$, during the updating process. If $\theta$ is randomly initialized as $\theta_{0}$, the initial value of $\mathcal{G}$, i.e., $\mathcal{G}_{0}=f_{\theta_{0}}(\mathcal{X}_{0},P)$ is far from the true $\mathcal{G}$. We find that the optimization process could exhibit turbulence and not produce the best results. Thus, an initialization problem is considered to produce a better initial value of $\theta$. We see that applying the blurring operator on both sides of $\mathcal{X}=\mathcal{G}\odot\hat{P}$ gives rise to $\mathcal{X}\otimes K\approx\mathcal{G}\odot(\hat{P}\otimes K)$ [22]. Thus, we propose to initialize $\theta$ by
$\theta_{0}^{*}=\arg\min_{\theta}\leavevmode\nobreak\ \|\hat{\mathcal{Y}}-f_{\theta}(\hat{\mathcal{Y}},P)\odot(\hat{P}\otimes K)\|,$ (13)
where $\hat{\mathcal{Y}}$ means the upsampled LRMS. Then, $\theta_{0}^{*}$ is used as the initial value of $\theta$ for iterative updating for the problem (7). In Algorithm 1, we summarize the entire process of implementing our PSDip. In addition, Fig. 2 shows the PSDip flow chart to take a comprehensive look at PSDip.
@@
0:  $\mathcal{X}_{0}=\hat{\mathcal{Y}}$, $\theta_{0}^{*}$ by (13)
@@ III-D Implementation details
We adopt PanNet [41] as the backbone of our network $f_{\theta}$. The network mainly contains convolutional layers and skip connections. In addition, we add ReLU activation to the final layer of PanNet to ensure that the output is always positive. It should be noted that the network structure is not exclusively specified. For the initialization problem (13), we use Adam to optimize $\theta$ until the objective converges, which takes about 8000 steps. The learning rate is set as $\mathrm{1e}^{-3}$. $\hat{P}$ is constructed by histogram-matching. That is each band of $\hat{P}$ is entirely translated and stretched such that the mean and standard values match those of the same band of LRMS. For the iterative update for the problem (7), we set $\alpha=2$ and $\beta=\mathrm{1e^{-3}}$. The iteration is also completed when the value of $\mathcal{L}(\mathcal{X},\theta)$ changes slowly, which takes about 3000 steps. That is, we set $T=3000$ in Algorithm 1.
@@ IV Experiments
PSDip 0.9088 $\pm$ 0.0227 0.0245 $\pm$ 0.0067 0.0684 $\pm$ 0.0179 349.2
In this section, we conduct several experiments to verify the effectiveness of the proposed PSDip. The following benchmark datasets are utilized, which can be downloaded from the website [^2] https://liangjiandeng.github.io/PanCollection.html. Specifically, we conduct reduced resolution experiments on WorldView-2 (WV2), WorldView-3 (WV3) and QuickBird (QB) test datasets. Each of them contains 20 sets of images, in which the HRMS, LRMS and PAN are included. The HRMS sizes of WV2, WV3 and QB are $256\times 256\times 8,256\times 256\times 8$ and $256\times 256\times 4$, respectively. The downscale factor $r$ is 4 for all datasets. That is the LRMS has the spatial size of $64\times 64$. We conduct full-resolution experiments on the QB test dataset, which also contains 20 sets of images but no HRMS is available. The full-resolution PAN in the QB dataset has the size of $512\times 512$ and the corresponding LRMS has the size of $128\times 128\times 4$. Besides, five images from the QB validation dataset are used for analyzing the trade-off parameter $\lambda$ and the setting of $f_{\theta}$. The original image value of each dataset ranges from 0 to 211. We normalize the data into [0,1] before the experiments.
@@ IV-A Pansharpening results
In Fig. 4-7, we visualize the pansharpening results of all methods on the test datasets. We select the $(8,2,1)$-th bands of WV2/WV3 data and the $(2,4,1)$-th bands of QB data to generate the pseudo-color images. The residual image is composed of the absolute values of the difference between the ground-truth HRMS and the restored HRMS. We see that the results of BDSD-PC, C-GSA, GLP-HPM, GLP-FS and LDPNet are a little over-smoothed. Particularly, in Fig. 5, The residual images of LRTCFPan and our PSDip obviously contain less image information. Besides, our method also tends to preserve more texture details than other methods, which reveals the effectiveness of the predicted $\mathcal{G}$ by our method.
@@ IV-B Analysis of $\mathcal{G}$ generated by PSDip
In the proposed PSDip, we estimate $\mathcal{G}$ by a neural network $f_{\theta}$. The network is first initialized by (13) and then iteratively updated together with $\mathcal{X}$ in (7). In Fig. 8, we illustrate the intermediate $\mathcal{G}$s generated from both processes to show how $\mathcal{G}$ changes. At the very beginning when $f_{\theta}$ has not been optimized at all, we see that the generated $\mathcal{G}_{0}$ already contains a few relevant image information. This observation confirms that the network structure can construct image structures even though the network parameters are randomly set. However, the image information of $\mathcal{G}_{0}$ does not meet our requirements for pan-sharpening. For the first 100 steps of the initialization process, the generated $\mathcal{G}$ quickly matches the “spectral” information, which is revealed by the “color change”. In the remaining steps of the initialization process, we see that $\mathcal{G}$ is gradually refined with spatial details as well as spectral information. These refinements make $\mathcal{G}$ look sharper. The final $\mathcal{G}$ of the initialization process looks more like the expected $\mathcal{G}$ but is of course still quite blurry. During the iterative updating process, we see that there is an apparent quality improvement on $\mathcal{G}$ compared with the initialization process. This is attributed to the formulated optimization problem that can more carefully establish the relationship between HRMS, LRMS and PAN. Besides, $\mathcal{G}$ is also incrementally incorporated with more details during this process.
@@
We can calculate three special intermediate $\mathcal{G}$s in the entire above optimization process. When initialization process (13) finishes, we get a initialized value of $\mathcal{G}$, i.e., $\mathcal{G}_{init}:=f_{\theta_{0}^{*}}(\hat{\mathcal{Y}},P)$. The final estimated $\mathcal{G}$ when the iterative updating process finishes is denoted as $\mathcal{G}_{T}:=f_{\theta_{T}}(\mathcal{X}_{T},P)$. Besides, we can also put $\mathcal{X}_{gt}$ into the final trained network $f_{\theta_{T}}$, the derived output is denoted as $\tilde{\mathcal{G}}:=f_{\theta_{T}}(\mathcal{X}_{gt},P)$. In Table V, we calculate the mean square error (MSE) between the ground-truth $\mathcal{G}=\mathcal{X}_{gt}\oslash\hat{P}$ and $\mathcal{G}_{init},\mathcal{G}_{T},\tilde{\mathcal{G}}$, respectively. The first line in Table V shows that the initialization (13) can produce rough values of $\mathcal{G}$ for the iterative updating. The second line shows that the final estimated $\mathcal{G}_{T}$ is more accurate than $\mathcal{G}_{init}$, which reveals the benefits of updating $\mathcal{G}$ in the proposed problem (7). Compared with the first two lines in Table V, the last line gets the smallest error. Specifically, this means that our trained $f_{\theta_{T}}$ could produce more accurate $\mathcal{G}$ when the inputs are “ideal”. Thus, the trained $f_{\theta}$ itself by our PSDip should exactly have learned how to predict $\mathcal{G}$ from $\mathcal{X}$ and $P$. Fig. 9 shows visualization of the ground-truth $\mathcal{G}$, our estimated $\mathcal{G}_{T}$ and $\tilde{\mathcal{G}}$. It can be seen that the coefficient $\mathcal{G}$ can be finely estimated by our PSDip.
@@
Figure 8: Visualization of $\mathcal{G}$ during the $f_{\theta}$ initialization process and iterative updating process for solving (7).
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@ IV-C Parameter analysis
There are mainly two parameters to be set. One is the only trade-off parameter $\lambda$ in the proposed problem (7). The other one is the step size $\alpha$ for updating $\mathcal{X}$ in (10). Five images from the QB validation dataset are used to analyze the two parameters. We present the results in Fig. 10. When $\lambda$ is less than 0.2, its growing value enables the generated coefficient $\mathcal{G}=f_{\theta}(\mathcal{X},P)$ to contribute more to the restored HRMS and the PSNR value then gradually rises. However, when $\lambda$ keeps growing exceeding 0.2, the spectral fidelity term $L_{p}$ occupies too much proportion in the optimization objective. The training of $f_{\theta}$ tends to be unstable because both $\mathcal{X}$ and $\theta$ are undetermined in $L_{p}$. Thus, we set $\lambda=0.1$ for all experiments to have a good balance between $L_{y}$ and $L_{p}$. For the step size $\alpha$, when its value is greater than 2, we see that the iterative updating also becomes unstable. Thus, we set $\alpha=2$ for all the experiments.
In Fig. 11, we plot the loss and PSNR trends in the initialization and iterative updating process on the QB test dataset. Let $\mathcal{L}_{init}$ denote the loss function in (13) for initializing $f_{\theta}$:
@@
The left image plots the averaged $\mathcal{L}_{init}$ on 20 images contained in the dataset. We see that the loss quickly drops at the first hundreds of steps, which is in consistent with the observation in Fig 8. The right image shows the trends of main objective $\mathcal{L}(\mathcal{X},\mathcal{X},P)$ in (7) and the corresponding PSNR values. The objective also quickly drops at the first dozens of steps and then gently converges. The PSNR value keeps gently growing when the iteration step is greater than 1000, in which stage we think the restored HRMS is gradually refined and added with more details. For all experiments, the initialization process takes 8000 optimization steps and the iterative updating process takes 3000 steps.
Figure 10: Analysis of parameters $\lambda$ and $\alpha$ using five images from QB validation dataset.
@@
Iterative updating
Figure 11: Trends of losses or PSNR in the initialization process and iterative updating process on QB test dataset.
@@ IV-D Ablation studies of $f_{\theta}$
In this section, we conduct several ablation studies about settings of the network $f_{\theta}$ of our PSDip. The experiments use five images from the QB validation dataset. Specifically, the following three cases are considered.
@@
case 2: $f_{\theta}$ still takes $\mathcal{X}$ and $P$ as inputs like PSDip. However, during the iterative updating process, only $\mathcal{X}$ is optimized and $\theta$ is fixed as $\theta_{0}^{*}$.
case 3: $f_{\theta}$ still takes $\mathcal{X}$ and $P$ as inputs like PSDip. However, it is not initialized as in (13) but directly optimized together with $\mathcal{X}$ in the iterative updating process.
The results are presented in Table VI. Comparing case 1 and our PSDip, we can see that our inputs $(\mathcal{X},P)$ could provide more information for generating $\mathcal{G}$ than random input $z$ and thus our PSDip has better performance. Results of case 2 show that updating $\theta$, which also means updating $\mathcal{G}$, for the proposed pansharpening problem (7) is useful, although the initialization process could produce a relatively “good” $\mathcal{G}$. Results of case 3 show that directly starting optimizing $f_{\theta}$ in the main objective (7) brings more uncertainty and does not produce satisfying results. Initializing $f_{\theta}$ using (13) is more efficient and could get better performance.
Table VI: Results of different settings of $f_{\theta}$ on QB validation dataset. The best results are in bold.

### v2

@@ III-C Optimization for the PSDip model
PSDip can be conveniently solved by alternating minimization [73][74]. Let $\mathcal{L}(\mathcal{X},\theta)$ denote the objective function of problem (8)
@@
Then, $\mathcal{X}$ can be simply updated by gradient descent:
@@
A good initial value of $\theta$ for the alternating minimization could help to stabilize the updating process and then help our model (6) to achieve better performance. We see that applying the blurring operator on both sides of $\mathcal{X}\approx\mathcal{G}\odot\hat{P}$ gives rise to $\mathcal{X}\otimes K\approx\mathcal{G}\odot(\hat{P}\otimes K)$ [25]. Thus, we initialize $\theta$ by
$\theta_{0}^{*}=\arg\min_{\theta}\leavevmode\nobreak\ \left\|\hat{\mathcal{Y}}-f_{\theta}(\hat{\mathcal{Y}},P)\odot(\hat{P}\otimes K)\right\|_{F},$ (14)
where $\hat{\mathcal{Y}}\approx\mathcal{X}\otimes K$ means the upsampled LRMS. $\hat{\mathcal{Y}}$ is also considered as the first input of $f_{\theta}$ since we do not have access to $\mathcal{X}$. Then, $\theta_{0}^{*}$ is used as the initial value of $\theta$ for alternating minimization for the problem (8). In Algorithm 1, we summarize the entire process of implementing our PSDip. Implementation details about Algorithm 1 are presented in Sec. III-D. In addition, Fig. 2 shows the PSDip flow chart to take a comprehensive look at PSDip.
@@
0:  $\mathcal{X}_{0}=\hat{\mathcal{Y}}$, $\theta_{0}^{*}$ by Eq. (14)
@@ III-D Implementation details
We adopt PanNet [45] as the backbone of our network $f_{\theta}$. The network mainly contains convolutional layers and skip connections. In addition, we add ReLU activation to the final layer of PanNet to ensure that the output is always positive. It should be noted that the network structure is not exclusively specified. In Sec. IV-G, we also present the results of two other networks. To construct $\hat{P}$, we first perform histogram-matching to generate $\hat{P}^{\prime}$, i.e., each band of $\hat{P}^{\prime}$ is entirely translated and stretched such that the mean and standard values match those of the same band of LRMS. To avoid zero-value in denominator, we then add a small value ($\varepsilon=1e-2$) to $\hat{P}^{\prime}$ and finally get $\hat{P}=\hat{P}^{\prime}+\varepsilon$. The upsampled LRMS $\hat{\mathcal{Y}}$ is derived by general bicubic interpolation of LRMS $\mathcal{Y}$. The kernel $K$ matches the MTF of multispectral sensors [72]. For the initialization problem (14), we use Adam to optimize $\theta$ until the objective converges, which takes about 8000 steps. The learning rate is constantly set as $\mathrm{1e}^{-3}$. For the alternating minimization for the main problem (8), we set $\alpha=2$ and $\beta=\mathrm{1e^{-3}}$ for all experiments. In each step of alternating minimization, $\mathcal{X}$ and $\theta$ are updated by one step of gradient descent and ADAM, respectively. The trade-off parameter $\lambda$ is set as 0.1 for all experiments. The iteration is also completed when the value of $\mathcal{L}(\mathcal{X},\theta)$ changes slowly, which takes about 3000 steps. That is, we set $T=3000$ in Algorithm 1.
@@ IV Experiments
PSDip 0.9102 $\pm$ 0.0238 0.0235 $\pm$ 0.0074 0.0680 $\pm$ 0.0184 349.2
In this section, we conduct several experiments to verify the effectiveness of the proposed PSDip. The following benchmark datasets are utilized, which can be downloaded from the website [^2] https://liangjiandeng.github.io/PanCollection.html. Specifically, we conduct reduced resolution experiments on WorldView-2 (WV2), WorldView-3 (WV3) and QuickBird (QB) test datasets. Each of them contains 20 sets of images, in which the HRMS, LRMS and PAN are included. The HRMS sizes of WV2, WV3 and QB are $256\times 256\times 8,256\times 256\times 8$ and $256\times 256\times 4$, respectively. The downscale factor $r$ is 4 for all datasets. That is the LRMS has the spatial size of $64\times 64$. We conduct full-resolution experiments on the QB test dataset, which also contains 20 sets of images but no HRMS is available. The full-resolution PAN in the QB dataset has the size of $512\times 512$ and the corresponding LRMS has the size of $128\times 128\times 4$. Besides, five images from the WV3 validation dataset are used for analyzing the trade-off parameter $\lambda$ and the setting of $f_{\theta}$. The original image value of each dataset ranges from 0 to 211. We normalize the data into [0,1] before the experiments.
@@ IV-A Pansharpening results
In Fig. 4-7, we visualize the pansharpening results of all methods on the test datasets. We select the $(8,2,1)$-th bands of WV2/WV3 data and the $(2,4,1)$-th bands of QB data to generate the pseudo-color images. The residual image is composed of the absolute difference between the pseudo-color HRMS and the pseudo-color restored HRMS. We see that the results of BDSD-PC, C-GSA, GLP-HPM, GLP-FS and LDPNet are a little over-smoothed. Particularly, in Fig. 5, The residual images of LRTCFPan and our PSDip obviously contain less image information. Besides, our method also tends to preserve more texture details than other methods, which reveals the effectiveness of the predicted $\mathcal{G}$ by our method.
@@ IV-B Analysis of $\mathcal{G}$ generated by PSDip
In the proposed PSDip, we estimate $\mathcal{G}$ by a neural network $f_{\theta}$. The network is first initialized by (14) and then iteratively updated together with $\mathcal{X}$ in (8). In Fig. 8, we illustrate the intermediate $\mathcal{G}$s generated from both processes to show how $\mathcal{G}$ changes. At the very beginning when $f_{\theta}$ has not been optimized at all, we see that the generated $\mathcal{G}_{0}$ already contains a few relevant image information. This observation confirms that the network structure can construct image structures even though the network parameters are randomly set. However, the image information of $\mathcal{G}_{0}$ does not meet our requirements for pan-sharpening. For the first 100 steps of the initialization process, the generated $\mathcal{G}$ quickly matches the “spectral” information, which is revealed by the “color change”. In the remaining steps of the initialization process, we see that $\mathcal{G}$ is gradually refined with spatial details as well as spectral information. These refinements make $\mathcal{G}$ look sharper. The final $\mathcal{G}$ of the initialization process looks more like the expected $\mathcal{G}$ but is of course still quite blurry. During the alternating minimization process, we see that there is an apparent quality improvement on $\mathcal{G}$ compared with the initialization process. This is attributed to the formulated optimization problem that can more carefully establish the relationship between HRMS, LRMS and PAN. Besides, $\mathcal{G}$ is also incrementally incorporated with more details during this process.
@@
We can calculate three special intermediate $\mathcal{G}$s in the entire above optimization process. When initialization process (14) finishes, we get a initialized value of $\mathcal{G}$, i.e., $\mathcal{G}_{init}:=f_{\theta_{0}^{*}}(\hat{\mathcal{Y}},P)$. The final estimated $\mathcal{G}$ when the alternating minimization process finishes is denoted as $\mathcal{G}_{T}:=f_{\theta_{T}}(\mathcal{X}_{T},P)$. Besides, we can also put $\mathcal{X}_{gt}$ into the final trained network $f_{\theta_{T}}$, the derived output is denoted as $\tilde{\mathcal{G}}:=f_{\theta_{T}}(\mathcal{X}_{gt},P)$. In Table V, we calculate the mean square error (MSE) between the ground-truth $\mathcal{G}=\mathcal{X}_{gt}\oslash\hat{P}$ and $\mathcal{G}_{init},\mathcal{G}_{T},\tilde{\mathcal{G}}$, respectively. The first line in Table V shows that the initialization (14) can produce rough values of $\mathcal{G}$ for the alternating minimization. The second line shows that the final estimated $\mathcal{G}_{T}$ is more accurate than $\mathcal{G}_{init}$, which reveals the benefits of updating $\mathcal{G}$ in the proposed problem (8). Compared with the first two lines in Table V, the last line gets the smallest error. Specifically, this means that our trained $f_{\theta_{T}}$ could produce more accurate $\mathcal{G}$ when the inputs are “ideal”. Thus, the trained $f_{\theta}$ itself by our PSDip should exactly have learned how to predict $\mathcal{G}$ from $\mathcal{X}$ and $P$. Fig. 9 shows visualization of the ground-truth $\mathcal{G}$, our estimated $\mathcal{G}_{T}$ and $\tilde{\mathcal{G}}$. It can be seen that the coefficient $\mathcal{G}$ can be finely estimated by our PSDip.
@@
Figure 8: Visualization of $\mathcal{G}$ during the $f_{\theta}$ initialization process and alternating minimization process for solving (8).
@@
Refer to caption MSE Refer to caption 0.0276 Refer to caption 0.0184
@@
Refer to caption MSE Refer to caption 0.0322 Refer to caption 0.0250
@@
Refer to caption MSE Refer to caption 0.0601 Refer to caption 0.0477
@@ IV-C Parameter analysis
There are mainly two parameters to be set. One is the only trade-off parameter $\lambda$ in the proposed problem (8). The other one is the step size $\alpha$ for updating $\mathcal{X}$ in (11). Five images from the WV3 validation dataset are used to analyze the two parameters. We present the results in Fig. 10. When $\lambda$ is less than 0.2, its growing value enables the generated coefficient $\mathcal{G}=f_{\theta}(\mathcal{X},P)$ to contribute more to the restored HRMS and the PSNR value then gradually rises. However, when $\lambda$ keeps growing exceeding 0.2, the spectral fidelity term $L_{p}$ occupies too much proportion in the optimization objective. The training of $f_{\theta}$ tends to be unstable because both $\mathcal{X}$ and $\theta$ are undetermined in $L_{p}$. Thus, we set $\lambda=0.1$ for all experiments to have a good balance between $L_{y}$ and $L_{p}$. For the step size $\alpha$, when its value is greater than 2, we see that the alternating minimization process also becomes unstable. Thus, we set $\alpha=2$ for all the experiments.
In Fig. 11, we plot the loss and PSNR trends in the initialization and alternating minimization process on the QB test dataset. Let $\mathcal{L}_{init}$ denote the loss function in (14) for initializing $f_{\theta}$:
@@
The left image plots the averaged $\mathcal{L}_{init}$ on 20 images contained in the dataset. We see that the loss quickly drops at the first hundreds of steps, which is in consistent with the observation in Fig 8. The right image shows the trends of main objective $\mathcal{L}(\mathcal{X},P)$ in (8) and the corresponding PSNR values. The objective also quickly drops at the first dozens of steps and then gently converges. The PSNR value keeps gently growing when the iteration step is greater than 1000, in which stage we think the restored HRMS is gradually refined and added with more details. For all experiments, the initialization process takes 8000 optimization steps and the alternating minimization process takes 3000 steps.
Figure 10: Analysis of parameters $\lambda$ and $\alpha$ using five images from WV3 validation dataset.
@@
Alternating minimization
Figure 11: Trends of losses or PSNR in the initialization and alternating minimization process on QB test dataset.
@@ IV-D Ablation studies of $f_{\theta}$
In this section, we conduct several ablation studies about settings of the network $f_{\theta}$ of our PSDip. The experiments use five images from the WV3 validation dataset. Specifically, the following three cases are considered.
@@
case 2: $f_{\theta}$ still takes $\mathcal{X}$ and $P$ as inputs like PSDip. However, during the alternating minimization process, only $\mathcal{X}$ is optimized and $\theta$ is fixed as $\theta_{0}^{*}$.
case 3: $f_{\theta}$ still takes $\mathcal{X}$ and $P$ as inputs like PSDip. However, it is not initialized by (14) but randomly initialized for the alternating minimization process.
The results are presented in Table VI. Comparing case 1 and our PSDip, we can see that our inputs $(\mathcal{X},P)$ could provide more information for generating $\mathcal{G}$ than random input $z$ and thus our PSDip has better performance. Results of case 2 show that updating $\theta$, which also means updating $\mathcal{G}$, for the proposed pansharpening problem (8) is useful, although the initialization process could produce a relatively “good” $\mathcal{G}$. Results of case 3 show that directly starting optimizing $f_{\theta}$ in the main objective (8) brings more uncertainty and does not produce satisfying results. Furthermore, in Fig. 12, we present the losses and PSNR values of case 3 and PSDip during the alternating minimization process. It shows that with a good initial value of $\theta$, the optimization process could be much more stable and also helps our model to get better performance.
Table VI: Results of different settings of $f_{\theta}$ on WV3 validation dataset. The best results are in bold.
@@
Initialization
Alternating minimization
Figure 12: Effectiveness of (14): trends of losses and PSNR in the alternating minimization process on WV3 validation dataset. Case3 does not initialize $f_{\theta}$ by $\min_{\theta}\mathcal{L}_{init}$ but uses a randomly initialized $f_{\theta}$ for alternating minimization.